40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 24 11:31:59 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling EX_fiber_MM.pdfEX_fiber_changes.png
    Reply  Tue Apr 24 18:54:15 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling 
       Reply  Wed Apr 25 19:09:37 2018, gautam, Configuration, ALS, New look EX Fiber coupling EX_fiber_MM.pdfIMG_6977.JPGIMG_6972.JPGIMG_6974.JPGIMG_6976.JPG
          Reply  Thu Apr 26 11:24:50 2018, gautam, Configuration, ALS, New look EX Fiber coupling - pol stability polStab.pdf
             Reply  Thu Apr 26 18:58:21 2018, Bruce, Configuration, ALS, New look EX Fiber coupling - pol stability 
             Reply  Fri Apr 27 01:36:02 2018, gautam, Configuration, ALS, IR ALS noise performance 2018_04_BeatMouth_POX.pdf2018_04_BeatMouth.pdfALSSpecgram.pdfALS_ASD.pdfALSstab.pdf
                Reply  Wed May 2 21:39:33 2018, gautam, Configuration, ALS, IR ALS for EY 
                   Reply  Fri May 4 21:17:57 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table 
                      Reply  Tue May 8 00:40:51 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table IMG_6986.JPGIMG_6987.JPG
                         Reply  Thu May 24 19:41:28 2018, gautam, Configuration, ALS, BeathMouth reinstalled on PSL table 
                      Reply  Thu May 24 13:06:17 2018, gautam, Configuration, ALS, DFD noises DFDcalib.pdfDFD_NB.pdf
                         Reply  Thu May 24 20:31:03 2018, gautam, Configuration, ALS, DFD noises TF_X_mag.pdfTF_X_phase.pdf
Message ID: 13807     Entry time: Wed May 2 21:39:33 2018     In reply to: 13796     Reply to this: 13817
Author: gautam 
Type: Configuration 
Category: ALS 
Subject: IR ALS for EY 

The new K6XS mounts I ordered have arrived. I want to install one of them at the Y-end. I can't find a picture of the current layout but it exists as there is a hardcopy affixed to the ETMY chamber door, Steve, can we dig this up and put it in the wiki? In any case, the current beam going into the fiber is the pickoff from the post-SHG harmonic separator. I'd like to change the layout a bit, and use a pickoff before the doubling oven, but looking at the optical table, this seems like a pretty involved task and would probably require large scale optical hardware rearrangement. In any case, the MM of the green beam into the Y-arm is <50%, so I would like to redo that as well. Does anyone know of a measurement of the mode from the Lightwave NPRO that is installed at EY? I think Annalisa is the one who installed this stuff, but I can't find an actual NPRO mode measurement in her elog thread.


Found it: elog4874, elog8436. I updated the laser inventory page to link the lasers in use to the most recent mode measurements I could find on the elog. I guess ideally we should also link the AM/PM response measurements.

------------------------------------------------------------------------------------------------------

SV  ETMY optical table layout  

     as of 3-31-2016

The oplev path was optimized with AR coated lenses and new He/Ne laser Jan 24, 2017

ELOG V3.1.3-